OpenStudy (anonymous):

A building lot in a city is shaped as a 30° -60° -90° triangle. The side opposite the 30° angle measures 41 feet.   a. Find the length of the side of the lot opposite the 60° angle b. Find the length of the hypotenuse of the triangular lot. c. Find the sine, cosine, and tangent of the 30° angle in the lot. Write your answers as decimals rounded to four decimal places.

OpenStudy (anonymous):

the hypotenuse is twice the length of the short side, so it is 82

OpenStudy (anonymous):

the long side is \(\sqrt{3}\) times the length of the short side, it is \(41\sqrt3\)

OpenStudy (anonymous):

@satellite73 lol..why do you leave me hanging so quickly..:(

OpenStudy (anonymous):

that was all right?

OpenStudy (anonymous):

\[\sin(30)=\frac{1}{2}, \cos(30)=\frac{\sqrt{3}}{2}, \tan(30)=\frac{1}{\sqrt{3}}\]

OpenStudy (anonymous):

you have to use a calculator to get the decimals

OpenStudy (anonymous):

oh ok thanks @satellite73! You are AWSOME! :)

OpenStudy (anonymous):

yw btw \(\cos(30)=.8660\) http://www.wolframalpha.com/input/?i=cos%2830%29